# Privacy-Preserving Retrieval-Augmented Generation with Differential Privacy

### v2

@@ Abstract
With the recent remarkable advancement of large language models (LLMs), there has been a growing interest in utilizing them in the domains with highly sensitive data that lies outside their training data. For this purpose, retrieval-augmented generation (RAG) is particularly effective—it assists LLMs by directly providing relevant information from the external knowledge sources. However, without extra privacy safeguards, RAG outputs risk leaking sensitive information from the external data source. In this work, we explore RAG under differential privacy (DP), a formal guarantee of data privacy. The main challenge with differentially private RAG is how to generate long accurate answers within a moderate privacy budget. We address this by proposing an algorithm that smartly spends privacy budget only for the tokens that require the sensitive information and uses the non-private LLM for other tokens. Our extensive empirical evaluations reveal that our algorithm outperforms the non-RAG baseline under a reasonable privacy budget of $\epsilon\approx 10$ across different models and datasets.
1 Introduction
@@
Large language models (LLMs) have shown a great deal of promise in a variety of applications. In particular, a major application of LLMs is in question-answering (QA). The practical adoption of these systems often involves domains whose data is highly sensitive. For instance, healthcare institutions might want to utilize their internal medical records to provide precise medical information and personal feedback, while legal firms can leverage their case archives to assist clients with legal research and documentation. One way to achieve such domain-specific QA is through retrieval-augmented generation (RAG) [5, 24, 17]. Here, we have a set of domain-specific documents; while answering a question, RAG retrieves a list of relevant documents and inputs them to LLMs as the context. However, even though this is effective for QA, RAG on a sensitive corpus can leak private information about individual documents in the corpus [40, 31, 4, 30]. This is particularly problematic when end users are outside the data-holding entity, e.g., patients interacting with a healthcare institution’s RAG system.
@@
Our goal in this paper is to prevent the information leakage of the sensitive external corpus by designing a privacy-preserving RAG system. For this purpose, we use differential privacy (DP) [10, 11] as a notion of privacy. Differential privacy guarantees privacy by ensuring that the participation of a single person’s data does not make much difference to the probability of any output. In our system, we assume that each RAG document comes from a single individual, and our goal is to ensure differential privacy on the eventual answer of the LLM.
@@
There are two aspects of the challenges with designing an effective RAG algorithm under DP. The first is how to fit differential privacy into the RAG framework, and the second is how to manage the privacy-utility tradeoffs. We address the first challenge by proposing an algorithm, DPVoteRAG, based on the sample-and-aggregate framework in DP [27]. Our algorithm prepares multiple LLM instances, or voters, feeds disjoint partitions of the sensitive corpus to them, and produces output tokens one by one each through the majority vote of the voters’ token outputs. Note, however, that LLMs often output many tokens in response to a question. This is detrimental to privacy—the composition property of differential privacy states that multiple calculations based on the same dataset lead to greater privacy degradation. To resolve this challenge, we design another algorithm, DPSparseVoteRAG, that spends a privacy budget only when we need to. More specifically, we take advantage of the fact in RAG that LLMs require the sensitive corpus only when generating tokens related to the knowledge. When not, outputs from LLMs without any context suffice. We formalize this idea with the sparse vector technique in DP [12, 13]—when voters agree with the non-private output of the LLM without contexts, we will simply output the non-private one without incurring a privacy budget. Consequently, our algorithm successfully generates sufficiently long, accurate responses under a reasonable privacy budget.
@@
2 Preliminaries & Problem Setting
2.1 Retrieval-Augmented Generation with Large Language Model
@@ 2.2 Differential Privacy
Definition 1 ($(\epsilon,\delta)$-Differential Privacy [10]).
@@
One of the key properties of DP is composition—sequential runs of differentially private algorithms also satisfy differential privacy. The composition property quantitatively captures the intuition that the more we release the information about the sensitive data, the worse the privacy guarantee becomes. More specifically, suppose $M_{1},\ldots,M_{T}$ are $(\epsilon_{0},\delta_{0})$-differentially private algorithms, which can be chosen adaptively based on previous outputs. Sequential composition theorem [10] states that the composed sequence of such algorithms guarantee $(T\epsilon_{0},T\delta_{0})$-differential privacy. Furthermore, advanced composition theorem [14, 20] states that the total privacy guarantee has $\epsilon=\mathcal{O}(\sqrt{T}\epsilon_{0})$.
2.2.1 Sparse Vector Technique
The sparse vector technique [12, 13] has originally emerged as the alternative of the composition in DP when we have such a large number of numerical queries that the composition theorem cannot provide a reasonable privacy guarantee but we are only interested in answers above some threshold. In such a case, the sparse vector technique algorithm, Sparse, reports whether each (noisy) query answer exceeds the threshold. It is shown that the privacy guarantee degrades by the number of queries above the threshold, instead of the total number of queries. Therefore, we save privacy budget by much when we expect only a few queries will be above the threshold. One can also report the (noisy) query outputs above the threshold by compromising privacy budget—the variant is called NumericSparse. The privacy guarantee follows by the composition of Sparse and differentially private query answering.
2.2.2 Differentially Private Generation via Sample-and-Aggregate
There has been a body of work on generating a token sequence by LLM with DP. The most common way is to borrow the idea of the sample-and-aggregate framework in DP [27]. To generate a single token, a set of LLMs, each depending on a disjoint subset of the sensitive dataset $D$, generates a token respectively. The generated tokens form an aggregate histogram of tokens, which is then carefully randomized with noise and only the most frequent token in the noisy histogram is published as the final output. The repetition of this process along with the composition theorem of DP yields the differentially private token sequence generation.
2.3 Problem Setting
@@
3 Differentially Private Retrieval-Augmented Generation with Sparse Vector Technique
Figure 2: Overview of DPSparseVoteRAG. It first retrieves the relevant documents to a question $x$ from sensitive external knowledge $D$ and partitions them uniform randomly (left). When generating $t$-th token (right), it takes a question $x$ and an interim answer up to $t-1$-th tokens, $y_{<t}$, as inputs and outputs a new token $y_{t}$. Components enclosed in red indicate that operations involve randomness for the privacy guarantee. DPVoteRAG works the same until computing the output token histogram $h_{t}$, but it immediately finds the majority token in $h_{t}$ afterwards.
@@ 3 Differentially Private Retrieval-Augmented Generation with Sparse Vector Technique
3.1 DPVoteRAG: Differentially Private Voting Algorithm for RAG
Algorithm 1 DPVoteRAG
@@ 3.1 DPVoteRAG: Differentially Private Voting Algorithm for RAG
1:  $T_{\max}\leftarrow$ maximum # of tokens to generate based on $(\epsilon_{\mathrm{token}},\delta_{\mathrm{token}})$ and $(\epsilon_{\mathrm{total}},\delta_{\mathrm{total}})$
2:  // Retrieval and random partition of the relevant documents
3:  $D_{x}\leftarrow$ Retrieve $mk$ most relevant documents $R(x,D;mk)$.
4:  $D_{x}^{1},\cdots,D_{x}^{m}\leftarrow$ Uniformly randomly partition $D_{x}$ into $m$ disjoint subsets.
@@
6:     // Generating the token histogram with the sensitive documents
@@
11:     // Producing a token from the histogram privately
@@
13:     // Halting if end of sequence
[…]
@@
18:  return $(y_{1},\ldots,y_{T_{\max}})$
@@
Given a prompt $x$ and external data source $D$, DPVoteRAG first retrieves $mk$ documents as $D_{x}$. Then it makes uniformly randomly partitions $D_{x}$ into $m$ disjoint datasets $D_{x}^{1},\ldots,D_{x}^{m}$ and each subset has exact size $k$. Then, for each $i=1,\ldots,m$, it feeds $k$ documents $D_{x}^{i}$ into the LLM along with the original prompt $x$, and generates a next token. It collects these tokens to form a histogram over the vocabulary. It remains to privately choose the most frequent element from the histogram. While it is generally hard to do so when the histogram dimension is large as in our setting, e.g., the vocabulary size of OPT [42] is 50272, there is a line of work in the community to overcome this difficulty. Following the work by [16], we integrate the LimitedDomain mechanism [9] into our algorithm. The mechanism enables us to reduce the histogram dimension significantly with some cost in a privacy budget and thus achieve a better privacy-utility tradeoff. By its design, the LimitedDomain mechanism possibly outputs the designed null token. In such a case, we halt the algorithm, or equivalently, regard that it outputs the end of sequence token.  [^2] We find that by choosing the appropriate size of reduced dimension, the LimitedDomain mechanism in our experiment rarely outputs the null token. Finally, we append the chosen token to the next input to the LLM. We repeat this process until we see the end of sequence token chosen or reach the maximum number of token generation, which is computed in advance from the per-token and total privacy budget  [^3] The maximum number of token generation is computed as follows. We first calculate the maximum numbers of composition with the sequential and advanced composition theorem [13] under the per-token privacy budget $(\epsilon_{\mathrm{token}},\delta_{\mathrm{token}})$ and total privacy budget $(\epsilon_{\mathrm{total}},\delta_{\mathrm{total}})$. Then, we take the maximum of two numbers of possible composition.. We present the concrete algorithm in Algorithm 1. The formal privacy analysis is as follows.
@@
The guarantee simply follows from the property of uniformly random partition, the privacy guarantee of the LimitedDomain mechanism and the composition theorem used to compute $T_{\max}$. The full proof is in Appendix A.
3.2 DPSparseVoteRAG: Differentially Private Voting Algorithm for RAG with Sparse Vector Technique
Algorithm 2 DPSparseVoteRAG
@@ 3.2 DPSparseVoteRAG: Differentially Private Voting Algorithm for RAG with Sparse Vector Technique
1:  // Privacy budget setup
2:  $(\epsilon_{\mathrm{token-RAG}},\delta_{\mathrm{token-RAG}})\leftarrow(\epsilon_{\mathrm{token}}/2,\delta_{\mathrm{token}})$, $\epsilon_{\mathrm{token-Lap}}\leftarrow\epsilon_{\mathrm{token}}/2$
3:  $c_{\max}\leftarrow$ maximum # of tokens to generate privately based on $(\epsilon_{\mathrm{token}},\delta_{\mathrm{token}})$ and $(\epsilon_{\mathrm{total}},\delta_{\mathrm{total}})$
4:  $c\leftarrow c_{\max}$, $\hat{\tau}\leftarrow\tau+\mathrm{Lap}(\nicefrac{{2}}{{\epsilon_{\mathrm{token-Lap}}}})$
5:  // Retrieval and random partition of the relevant documents
6:  $D_{x}\leftarrow$ Retrieve $mk$ most relevant documents $R(x,D;mk)$.
7:  $D_{x}^{1},\cdots,D_{x}^{m}\leftarrow$ Uniformly randomly partition $D_{x}$ into $m$ disjoint subsets.
@@
9:     // Generating the non-private token and token histogram with the sensitive documents
@@
15:     // Producing a token from the histogram privately only when $y_{t}^{\mathrm{non-RAG}}$ is uncommon in $h_{t}$
@@
19:        // The privacy budget is only consumed when $y_{t}$ is from the histogram
@@
24:     // Halting if end of sequence or the privacy budget has been exhausted
[…]
@@
29:  return $(y_{1},\ldots,y_{T_{\max}})$
@@
We address this by incorporating the sparse vector technique into our voting algorithm, yielding our improved algorithm DPSparseVoteRAG. In particular, before we apply private voting among generated tokens, we check if the generated tokens coincide with the token generated by the LLM without retrieved documents appended, i.e., the input is composed of the prompt and previously generated tokens only. We continue to the voting only when they do not coincide. Otherwise, we use the LLM output without retrieved documents. It is shown from the analysis of the sparse vector technique that the consumed privacy budget scales with the number of times that it uses the private voting, not with the total number of generated tokens. Consequently, the resulting algorithm, shown in Algorithm 2, enables us to spend a privacy budget only when it needs sensitive information. We note that as a result of this change in our algorithm, we compute the maximum number of tokens to generate with private voting, $c_{\max}$, from the per-token privacy budget $(\epsilon_{\mathrm{token}},\delta_{\mathrm{token}})$ and total privacy budget $(\epsilon_{\mathrm{total}},\delta_{\mathrm{total}})$, instead of the maximum number of token generation, $T_{\max}$, as in DPVoteRAG but in the same way to compute $T_{\max}$. The formal privacy analysis of this algorithm is as follows. The guarantee holds due to the privacy guarantee of the LimitedDomain mechanism and the NumericSparse algorithm [13].
@@
4 Experiment
We investigate how our differentially private voting RAG algorithms (Algorithms 1 and 2) work. Specifically, we ask the following questions:

### v3

@@ Abstract
With the recent remarkable advancement of large language models (LLMs), there has been a growing interest in utilizing them in the domains with highly sensitive data that lies outside their training data. For this purpose, retrieval-augmented generation (RAG) is particularly effective—it assists LLMs by directly providing relevant information from the external knowledge sources. However, without extra privacy safeguards, RAG outputs risk leaking sensitive information from the external data source. In this work, we explore RAG under differential privacy (DP), a formal guarantee of data privacy. The main challenge with differentially private RAG is how to generate long accurate answers within a moderate privacy budget. We address this by proposing an algorithm that smartly spends privacy budget only for the tokens that require the sensitive information and uses the non-private LLM for other tokens. Our extensive empirical evaluations reveal that our algorithm outperforms the non-RAG baseline under a reasonable privacy budget of $\epsilon\approx 10$ across different models and datasets. The code for reproducing our results is at https://github.com/tacchan7412/DPRAG.
I Introduction
@@
Large language models (LLMs) have shown a great deal of promise in a variety of applications. In particular, a major application of LLMs is in question-answering (QA). The practical adoption of these systems often involves domains whose data is highly sensitive. For instance, healthcare institutions might want to utilize their internal medical records to provide precise medical information and personal feedback, while legal firms can leverage their case archives to assist clients with legal research and documentation. One way to achieve such domain-specific QA is through retrieval-augmented generation (RAG) [chen_reading_2017, lewis_retrieval-augmented_2020, hsia2024ragged]. Here, we have a set of domain-specific documents; while answering a question, RAG retrieves a list of relevant documents and inputs them to LLMs as the context. However, even though this is effective for QA, RAG on a sensitive corpus can leak private information about individual documents in the corpus [zeng_good_2024-2, qi_follow_2024, chaudhari_phantom_2024, peng_data_2024]. This is particularly problematic when end users are outside the data-holding entity, e.g., patients interacting with a healthcare institution’s RAG system.
@@
Our goal in this paper is to prevent the information leakage of the sensitive external corpus by designing a privacy-preserving RAG system. For this purpose, we use differential privacy (DP) [dwork_our_2006-1, dwork_calibrating_2006] as a notion of privacy. Differential privacy guarantees privacy by ensuring that the participation of a single person’s data does not make much difference to the probability of any output. In our system, we assume that each RAG document comes from a single individual, and our goal is to ensure differential privacy on the eventual answer of the LLM.
@@
There are two aspects of the challenges with designing an effective RAG algorithm under DP. The first is how to fit differential privacy into the RAG framework, and the second is how to manage the privacy-utility tradeoffs. We address the first challenge by proposing an algorithm, DPVoteRAG, based on the sample-and-aggregate framework in DP [nissim_smooth_2007]. Our algorithm prepares multiple LLM instances, or voters, feeds disjoint partitions of the sensitive corpus to them, and produces output tokens one by one each through the majority vote of the voters’ token outputs. Note, however, that LLMs often output many tokens in response to a question. This is detrimental to privacy—the composition property of differential privacy states that multiple calculations based on the same dataset lead to greater privacy degradation. To resolve this challenge, we design another algorithm, DPSparseVoteRAG, that spends a privacy budget only when we need to. More specifically, we take advantage of the fact in RAG that LLMs require the sensitive corpus only when generating tokens related to the knowledge. When not, outputs from LLMs without any context suffice. We formalize this idea with the sparse vector technique in DP [dwork_complexity_2009, dwork_algorithmic_2014]—when voters agree with the non-private output of the LLM without contexts, we will simply output the non-private one without incurring a privacy budget. Consequently, our algorithm successfully generates sufficiently long, accurate responses under a reasonable privacy budget.
@@
II Preliminaries & Problem Setting
II-A Retrieval-Augmented Generation with Large Language Model
@@ II-B Differential Privacy
Definition 1 ($(\epsilon,\delta)$-Differential Privacy [dwork_our_2006-1]).
@@
One of the key properties of DP is composition—sequential runs of differentially private algorithms also satisfy differential privacy. The composition property quantitatively captures the intuition that the more we release the information about the sensitive data, the worse the privacy guarantee becomes. More specifically, suppose $M_{1},\ldots,M_{T}$ are $(\epsilon_{0},\delta_{0})$-differentially private algorithms, which can be chosen adaptively based on previous outputs. Sequential composition theorem [dwork_our_2006-1] states that the composed sequence of such algorithms guarantee $(T\epsilon_{0},T\delta_{0})$-differential privacy. Furthermore, advanced composition theorem [dwork_boosting_2010-1, kairouz_composition_2017-1] states that the total privacy guarantee has $\epsilon=\mathcal{O}(\sqrt{T}\epsilon_{0})$.
II-B1 Sparse Vector Technique
The sparse vector technique [dwork_complexity_2009, dwork_algorithmic_2014] has originally emerged as the alternative of the composition in DP when we have such a large number of numerical queries that the composition theorem cannot provide a reasonable privacy guarantee but we are only interested in answers above some threshold. In such a case, the sparse vector technique algorithm, Sparse, reports whether each (noisy) query answer exceeds the threshold. It is shown that the privacy guarantee degrades by the number of queries above the threshold, instead of the total number of queries. Therefore, we save privacy budget by much when we expect only a few queries will be above the threshold.
We state the AboveThreshold in Algorithm 1 and state their guarantee as below.
0: A private database D, an adaptively chosen stream of sensitivity 1 queries $f_{1}$, $\cdots$, and a threshold $\tau$.
0: A stream of responses $a_{1}$, $\cdots$.
1: Let $\hat{\tau}=\tau+\mathrm{Lap}\left(\frac{2}{\varepsilon}\right)$.
2:for Each query $i$ do
3:  Let $v_{i}=\mathrm{Lap}\left(\frac{4}{\varepsilon}\right)$
4:  if $f_{i}(D)+v_{i}\geq\hat{\tau}$ then
5:   Output $a_{i}=1$, Halt.
6:  else
7:   Output $a_{i}=0$.
8:  end if
9:end for
Algorithm 1 AboveThreshold [dwork_algorithmic_2014]
Theorem 1.
ALgorithm 1 is $(\varepsilon,0)$-DP.
II-B2 Differentially Private Generation via Sample-and-Aggregate
There has been a body of work on generating a token sequence by LLM with DP. The most common way is to borrow the idea of the sample-and-aggregate framework in DP [nissim_smooth_2007]. To generate a single token, a set of LLMs, each depending on a disjoint subset of the sensitive dataset $D$, generates a token respectively. The generated tokens form an aggregate histogram of tokens, which is then carefully randomized with noise and only the most frequent token in the noisy histogram is published as the final output. The repetition of this process along with the composition theorem of DP yields the differentially private token sequence generation.
II-C Problem Setting
@@
III Differentially Private Retrieval-Augmented Generation with Sparse Vector Technique
Figure 2: Overview of DPSparseVoteRAG. It first retrieves the relevant documents to a question $x$ from sensitive external knowledge $D$ and partitions them uniform randomly (left). When generating $t$-th token (right), it takes a question $x$ and an interim answer up to t1t-1-th tokens, $y_{<t}$, as inputs and outputs a new token $y_{t}$. Components enclosed in red indicate that operations involve randomness for the privacy guarantee. DPVoteRAG works the same until computing the output token histogram $h_{t}$, but it immediately finds the majority token in $h_{t}$ afterwards.
@@ III Differentially Private Retrieval-Augmented Generation with Sparse Vector Technique
III-A DPVoteRAG: Differentially Private Voting Algorithm for RAG
@@
[…]
2:{Retrieval and random partition of the relevant documents}
[…]
@@
6:  {Generating the token histogram with the sensitive documents}
@@
11:  {Producing a token from the histogram privately}
@@
13:  {Halting if end of sequence}
14:  if $y_{t}=\texttt{<EOS>}$ then
@@
118:return $(y_{1},\ldots,y_{T_{\max}})$
Algorithm 2 DPVoteRAG
@@
Given a prompt $x$ and external data source $D$, DPVoteRAG first retrieves $mk$ documents as $D_{x}$. Then it makes uniformly randomly partitions $D_{x}$ into $m$ disjoint datasets $D_{x}^{1},\ldots,D_{x}^{m}$ and each subset has exact size $k$. Then, for each $i=1,\ldots,m$, it feeds $k$ documents $D_{x}^{i}$ into the LLM along with the original prompt $x$, and generates a next token. It collects these tokens to form a histogram over the vocabulary. It remains to privately choose the most frequent element from the histogram. While it is generally hard to do so when the histogram dimension is large as in our setting, e.g., the vocabulary size of OPT [zhang_opt_2022-1] is 50272, there is a line of work in the community to overcome this difficulty. Following the work by hong2024dpopt, we integrate the LimitedDomain mechanism [durfee_practical_2019-1] into our algorithm. The mechanism enables us to reduce the histogram dimension significantly with some cost in a privacy budget and thus achieve a better privacy-utility tradeoff. By its design, the LimitedDomain mechanism possibly outputs the designed null token. In such a case, we halt the algorithm, or equivalently, regard that it outputs the end of sequence token.  [^3] We find that by choosing the appropriate size of reduced dimension, the LimitedDomain mechanism in our experiment rarely outputs the null token. Finally, we append the chosen token to the next input to the LLM. We repeat this process until we see the end of sequence token chosen or reach the maximum number of token generation, which is computed in advance from the per-token and total privacy budget  [^4] The maximum number of token generation is computed as follows. We first calculate the maximum numbers of composition with the sequential and advanced composition theorem [dwork_algorithmic_2014] under the per-token privacy budget $(\epsilon_{\mathrm{token}},\delta_{\mathrm{token}})$ and total privacy budget $(\epsilon_{\mathrm{total}},\delta_{\mathrm{total}})$. Then, we take the maximum of two numbers of possible composition.. We present the concrete algorithm in Algorithm 2. The formal privacy analysis is as follows.
@@
The guarantee simply follows from the property of uniformly random partition, the privacy guarantee of the LimitedDomain mechanism and the composition theorem used to compute $T_{\max}$.
Proof.
For any two neighboring datasets $D$ and $D^{\prime}$, the retrieved datasets are $D_{x}=R(x,D;mk)$ and $D_{x}^{\prime}=R(x,D^{\prime};mk)$ and we denote $L_{x}=(d_{1},\cdots,d_{mk})$ and $L_{x}^{\prime}=(d_{1}^{\prime},\cdots,d_{mk}^{\prime})$. We only need to show for any set of outcomes $S$, $\Pr_{A,\pi}[A(L_{x}^{\pi})\in S]\leq\exp(\epsilon)\Pr_{A,\pi}[A((L_{x}^{\prime})^{\pi})\in S]+\delta.$ First of all, $D_{x}$ and $D_{x}^{\prime}$ have at most one different document (without considering the order). Therefore we can define another list of documents $L^{\prime\prime}_{x}=(d_{1}^{\prime\prime},\cdots,d_{mk}^{\prime\prime})$, such that $L^{\prime\prime}_{x}$ is some ranking of $D_{x}$’ and it differs at most one position from $L$. Notice that $L^{\prime}_{x}$ and $L^{\prime\prime}_{x}$ are only different at orders. Therefore $(L_{x}^{\prime})^{\pi}$ and $(L_{x}^{\prime\prime})^{\pi}$ have same distributions and as a consequence $A((L_{x}^{\prime})^{\pi})$ and $A((L_{x}^{\prime\prime})^{\pi})$ have same distributions.
III-B DPSparseVoteRAG: Differentially Private Voting Algorithm for RAG with Sparse Vector Technique
@@
1:{Privacy budget setup}
[…]
5:{Retrieval and random partition of the relevant documents}
[…]
@@
9:  {Generating the non-private token and token histogram with the sensitive documents}
@@
15:  {Producing a token from the histogram privately only when $y_{t}^{\mathrm{non-RAG}}$ is uncommon in $h_{t}$}
@@
19:   {The privacy budget is only consumed when $y_{t}$ is from the histogram}
@@
24:  {Halting if end of sequence or the privacy budget has been exhausted}
25:  if $y_{t}=\texttt{<EOS>}$ or $c=0$ then
@@
129:return $(y_{1},\ldots,y_{T_{\max}})$
Algorithm 3 DPSparseVoteRAG
@@
We address this by incorporating the sparse vector technique into our voting algorithm, yielding our improved algorithm DPSparseVoteRAG. In particular, before we apply private voting among generated tokens, we check if the generated tokens coincide with the token generated by the LLM without retrieved documents appended, i.e., the input is composed of the prompt and previously generated tokens only. We continue to the voting only when they do not coincide. Otherwise, we use the LLM output without retrieved documents. It is shown from the analysis of the sparse vector technique that the consumed privacy budget scales with the number of times that it uses the private voting, not with the total number of generated tokens. Consequently, the resulting algorithm, shown in Algorithm 3, enables us to spend a privacy budget only when it needs sensitive information. We note that as a result of this change in our algorithm, we compute the maximum number of tokens to generate with private voting, $c_{\max}$, from the per-token privacy budget $(\epsilon_{\mathrm{token}},\delta_{\mathrm{token}})$ and total privacy budget $(\epsilon_{\mathrm{total}},\delta_{\mathrm{total}})$, instead of the maximum number of token generation, $T_{\max}$, as in DPVoteRAG but in the same way to compute $T_{\max}$. The formal privacy analysis of this algorithm is as follows. The guarantee holds due to the privacy guarantee of the LimitedDomain mechanism and the AboveThreshold algorithm [dwork_algorithmic_2014].
@@
Proof.
Similar to the proof of Theorem 2, we only need to prove that if $\mathcal{D}_{x}=(D_{x}^{1},\cdots,D_{x}^{m})$ at step 7 and $\mathcal{D}_{x}^{\prime}=(D_{x}^{1^{\prime}},\cdots,D_{x}^{m^{\prime}})$ differ at most one set $D_{x}^{i}$, $\Pr(y_{1},\cdots,y_{T_{\max}})\leq\varepsilon\cdot\Pr(y_{1}^{\prime},\cdots,y_{T_{\max}}^{\prime})+\delta$.
Denote $t_{\gamma}$ as the first time step that holds $c=\gamma$ at the beginning of this time step, e.g. $t_{c_{\max}}=1$. We can first prove that $\Pr(y_{t_{\gamma}\leq t<t_{\gamma-1}}|\mathcal{D}_{x},y_{t<t_{\gamma}})\leq\varepsilon_{\rm token}\cdot\Pr(y_{t_{\gamma}^{\prime}\leq t<t_{\gamma-1}^{\prime}}^{\prime}|\mathcal{D}_{x}^{\prime},y_{t<t_{\gamma}^{\prime}}^{\prime})+\delta_{\rm token}$ if (1) $t_{\gamma}=t_{\gamma}^{\prime}$ and (2) $y_{t<t_{\gamma}}=y_{t<t_{\gamma}}^{\prime}$.
This can be proved by two parts. First, because $t_{\gamma}=t_{\gamma}^{\prime}$,
$\Pr(t_{\gamma-1}|\mathcal{D}_{x},y_{t<t_{\gamma}})\leq\varepsilon_{\rm token}/2\cdot\Pr(t_{\gamma-1}|\mathcal{D}_{x}^{\prime},y_{t<t_{\gamma}^{\prime}}^{\prime}).$
This is implied by applying the sparse vector technique presented in Algorithm 1 and Theorem 1 to analyze our algorithm (step 17-20). Furthermore, if $t_{\gamma-1}=t_{\gamma-1}^{\prime}$,
$\Pr(y_{t_{\gamma}\leq t<t_{\gamma-1}}|\mathcal{D}_{x},y_{t<t_{\gamma}})=\Pr(y_{t_{\gamma}\leq t<t_{\gamma-1}}^{non-RAG},y_{t_{\gamma-1}}^{DP}|\mathcal{D}_{x},y_{t<t_{\gamma}})$
$\Pr(y_{t_{\gamma}^{\prime}\leq t<t_{\gamma-1}^{\prime}}^{non-RAG},y_{t_{\gamma-1}^{\prime}}^{DP}|\mathcal{D}_{x},y_{t<t_{\gamma}^{\prime}}^{\prime})$
$\leq\varepsilon_{\rm token}/2\cdot\Pr(y_{t_{\gamma}^{\prime}\leq t<t_{\gamma-1}^{\prime}}^{non-RAG},y_{t_{\gamma-1}}^{DP^{\prime}}|\mathcal{D}_{x}^{\prime},y_{t<t_{\gamma}^{\prime}}^{\prime})+\delta_{\rm token}/2$
$\Pr(y_{t_{\gamma}^{\prime}\leq t<t_{\gamma-1}^{\prime}}^{\prime}|\mathcal{D}_{x}^{\prime},y_{t<t_{\gamma}}^{\prime})$
where the first and the last equality come from the definition of the algorithm (step 17-23), the second equality holds because we assume $t_{\gamma}=t_{\gamma}^{\prime}$, $y_{t<t_{\gamma}}=y_{t<t_{\gamma}}^{\prime}$ and $t_{\gamma-1}=t_{\gamma-1}^{\prime}$, and the inequality comes from the DP guarantee by the LimitedDomain mechanism.
Lastly, our algorithm must stop before $c=0$, means that our algorithm is a composition of at most $c_{\max}$ steps of $(\varepsilon_{\rm token},\delta_{\rm token})$-DP. As shown in step 3 in our algorithm, $c_{\max}$ is picked to guarantee that the composition of $c_{\max}$ steps of $(\varepsilon_{\rm token},\delta_{\rm token})$-DP is $(\varepsilon_{\rm total},\delta_{\rm total})$-DP. Therefore, our algorithm is $(\varepsilon_{\rm total},\delta_{\rm total})$-DP. ∎
IV Experiment
We investigate how our differentially private voting RAG algorithms (Algorithms 2 and 3) work. Specifically, we ask the following questions: